# Biologically-Inspired Technologies: Integrating Brain-Computer Interface and Neuromorphic Computing for Human Digital Twins

### v2

@@ Abstract
Index Terms:
Human Digital Twin, Human-Centric Systems, Brain-Computer Interface, Spiking Neural Networks.
@@ 2. Q2)
    Given the need for continuous updating of HDT to adapt to real-time changes of user, what strategies can be developed to reduce the energy consumption involved in these data-intensive operations?
@@ 3. Q3)
  • We introduce a novel Bio-inspired HDT (Bio-HDT) framework that utilizes Brain-Computer Interface (BCI) sensor technology to collect human brain signals as the primary data source for constructing HDTs. This framework harnesses brain signals to capture real-time behavioral, psychological, physiological, and observational information of individuals, thereby providing richer data and a holistic view for constructing immersive and personalized HDTs. This approach not only enhances the precision and fidelity of the HDTs but also significantly reduces the reliance on heterogeneous IoT devices, simplifying both data collection and processing. Consequently, it ensures high QoS provided by MSPs and QoE perceived by users (For Q1).
  • We propose a Bio-inspired neuromorphic computing learning model based on the Spiking Neural Networks (SNNs) for efficiently processing extracted brain signals. Unlike traditional Artificial Neural Networks (ANNs) that process information continuously and require constant computational activity, SNNs utilize discrete, brain-like spikes to transmit information between neurons. This method more accurately reflects the natural communication processes of brain neurons, thereby enhancing data processing efficiency and reducing the energy consumption specifically involved in the real-time updating of HDTs (For Q2).
@@ II-A Human-centric Metaverse and Human Digital Twin Framework
As the next generation of the Internet, the human-centric Metaverse extends beyond providing a virtual space for gaming. It is envisioned as a new paradigm that leverages users’ behavioral, psychological, physiological, and observational data to bridge the virtual and physical worlds, thereby profoundly transforming human lifestyles. HDTs, similar to the critical role of digital twins (DT) in manufacturing [2], play a vital role in improving the human-centric Metaverse. For instance, they allow simulation and prediction of physiological and psychological states of humans, facilitating personalized medical treatments [1] and health management programs [3]. Besides, HDTs are also able to assist in generating intelligent avatars, creating highly customized environments that mimic real-world interactions and scenarios, thus offering unparalleled opportunities for users.
As shown in Fig. 1, the current HDT framework operates cyclically by executing the following steps: (1) Sensors embedded in IoT devices continuously collect user real-time behavioral data; (2) The data are then transmitted to the MSP; (3) The MSP processes and analyzes the data using advanced software and algorithms. After analyzing these data, MSPs are able to generate the user’s HDT. Currently, this framework has been utilized primarily in healthcare. For example, HDTs are integrated with wearable devices such as smartwatches, fitness trackers, and health sensors to continuously monitor vital factors like heart rate, blood pressure, and glucose levels. This data can be synchronized with medical records and utilized to generate real-time insights into an individual’s health, enabling early diagnosis of potential issues and the optimization of personalized healthcare strategies [1, 3, 4].
@@ II-B Challenges of the Current Human Digital Twins
Figure 1: The current HDT framework. Wearable devices collect observational, physiological, and behavioral data. These data are transmitted in real-time to the MSP through wireless links. The MSP then employs a range of customized software and models to construct and continuously update the HDT.
@@ II-B2 Energy Consumption (Q2)
To efficiently analyze these collected data, advanced models and algorithms based on deep learning have been deployed. However, they are primarily driven by ANNs, which are resource-intensive and require substantial computational power for both training and inference phases [5]. This intense energy demand presents significant challenges for these IoT devices with limited energy. First, the high power requirement for real-time data processing and analysis can lead to frequent battery depletion, disrupting continuous monitoring and data collection essential for maintaining accurate and dynamic HDTs. Moreover, the strain on device endurance due to persistent high-load operations can degrade device performance over time, increasing maintenance costs and potentially compromising the quality of user interaction within the Metaverse.
@@ III Brain-Computer Interface-Enhanced Human Digital Twin
In this section, we illustrate BCI technology and its current applications across various domains. We then propose a novel Bio-HDT framework that leverages advanced BCI technology, focusing on the potential for improved customization, efficiency, and interaction within the human-centric Metaverse. Finally, we discuss the unique challenges involved in deploying BCIs within HDT frameworks.
Figure 2: An illustration of the proposed framework for the biologically-inspired human digital twin (Bio-HDT). This process begins with the extraction of brain signals by a Brain-Computer Interface-driven Head-mounted Display (BCI-driven HMD). The signals are analyzed using the Spiking Neural Network to identify human behavior and states such as psychological state like stress and physiological state like fatigue, which are then mapped into the HDT. This framework enables MSPs to deliver personalized services tailored to the user’s real-time activities within a human-centric Metaverse.
@@ III-A Brain Signal and Brain-Computer Interface
Unlike previous works [1, 3, 4], this work leverages brain signals as a novel data source for constructing HDTs. Specifically, brain signals offer a multifaceted view into human nature, crucial for modeling the human activities. They contain critical behavioral, psychological, physiological, and observational information, enabling the decoding of subconscious and conscious responses to environmental stimuli. These insights serve as windows into the internal processes that govern human actions, perceptions, and interactions, presenting their significant potential for constructing HDTs. By harnessing these signals, we can develop a human-centric Metaverse where each user’s HDT is authentically driven by their unique brain activities, ensuring a more immersive and personalized virtual experience.
To this end, BCI technology that connects the brain signals and physical devices has been considered. As a direct communication pathway between the human brain and externals machines, devices, and system, BCI operates by monitoring brain activities and collecting brain signals through sensors attached to the human scalp (e.g., non-invasive BCI [6]), processing the data to detect specific patterns related to intended actions, and translating these patterns into actionable commands. This allows users to control devices or perform tasks without physical movement [7]. It has been successfully applied in various domains, demonstrating its broad potential and utility.
  • Medical: BCIs provide individuals with spinal cord injuries or neurodegenerative diseases greater autonomy by enabling control over prosthetic limbs, communication systems, and other assistive technologies. Additionally, BCIs are used in clinical settings to assess brain injuries and diagnose brain death, aiding in critical decision-making processes in healthcare [7].
  • Mental Health: BCIs are transforming the diagnosis and treatment of mental health disorders, including conditions like epilepsy, strokes, and sleep disorders. Their ability to detect subtle neurological signals enables them invaluable in managing neurodegenerative diseases such as Parkinson’s [8] and Alzheimer’s [9].
  • Entertainment: BCIs are revolutionizing gaming and immersive experiences by enabling users to control avatars and interact with virtual environments using only their thoughts. This technology significantly enhances user immersion and interaction, setting the stage for more intuitive and responsive VR systems [6, 10].
@@ III-B Biologically-inspired Human Digital Twins
As analyzed above, brain signals encompass underlying information about human activities. These activities provide a holistic view of an individual’s real-time states across multiple dimensions, which is particularly useful in contexts such as healthcare, personalized technology interactions, and immersive Metaverse application. By capturing and analyzing these various human behavior and states information, BCI technology assists in creating more personalized and responsive systems that adapt to individual needs and changes in their condition. Therefore, we propose the Bio-HDT framework that utilizes brain signals as the data source for conducting HDT and incorporates BCI technology to efficiently capture these signals. As illustrated in Fig. 2, the BCI-driven Head-mounted Display (HMD) system within the Bio-HDT framework continuously captures brain signals through embedded BCI sensors, which are then analyzed using a neuromorphic computing learning model that mimics brain function through SNNs. After real-time extraction and analysis of these brain signals, the Bio-HDT is constructed and synchronized with the corresponding human states in the physical world. Utilizing this real-time Bio-HDT, MSPs can identify and even predict fluctuations in a user’s psychology and physical intentions. Consequently, this framework is able MSPs to generate personalized services for the users by monitoring their brain signals (for Q1).
@@
Specifically, the Bio-HDT operates as a two-layer bio-inspired scheme. The first layer bio-inspired captures brain signals, which reflect the user’s cognitive, emotional, and behavioral states, providing an authentic representation of the brain’s biological processes. These brain signals are continuously monitored and collected through embedded BCI sensors in the HMD system, ensuring real-time feedback from the user. The second layer bio-inspired processes these brain signals using SNNs, a neuromorphic computing scheme that mimics the way neurons communicate in the brain. Different from traditional neural networks, SNNs process data in a manner that mirrors the brain’s natural signaling mechanisms, interpreting neural patterns through the timing and frequency of spikes, which is detailed in Section IV. This allows MSPs to accurately extract, analyze, and respond to users’ brain signals, generating personalized content in real time.
@@
By efficiently combining the BCI technology, the Bio-HDT framework offers several advantages. First, it minimizes the reliance on multiple IoT devices and external sensors traditionally required to monitor a user’s physical state. Conventional HDT systems rely on complex sensor networks, resulting in hardware demands, synchronization challenges, and increased energy consumption. By directly leveraging brain signals through the BCI, the Bio-HDT framework gathers critical multifaceted data (e.g., cognition, emotion, and behavior) without relying on an array of external sensors, thereby eliminating inefficiencies, reducing the investment in devices, and lowering computational load and data breach risk. Moreover, these brain signals are extracted unconsciously, without requiring explicit physical actions, preserving the user’s immersive experience and providing a more nuanced representation of their state. This facilitates more natural and immersive interactions within the Metaverse.
@@
Therefore, the Bio-HDT framework not only addresses the challenge of customizing HDTs for individual users (for Q1) but also minimizes energy consumption (for Q2) and reduces the risk of data breaches (for Q3) by eliminating the need for multiple data collection devices. Additionally, the real-time interaction between brain signals and the decision-making systems enhances the accuracy, responsiveness, and immersiveness of HDTs. This new paradigm brings users closer to a fully realized human-centric Metaverse. It is worth noting that the proposed Bio-HDT system can also serve humans effectively beyond the Metaverse environment. In such cases, the Bio-HDT system can be used to monitor various aspects of human states [7].
III-C Challenges
Although Bio-HDT framework that utilizes brain signals and BCI technology offers rich information for updating HDTs and unprecedentedly supporting the development of the human-centric Metaverse, several concerns must be addressed to fully realize its potential.
III-C1 Efficient Processing of Brain Signals
Brain signals are inherently temporal and dynamic, necessitating that deep learning models process large volumes of data to decipher fluctuating patterns and rhythms of brain activity. This complexity significantly complicates real-time interpretation. Traditional ANNs, which typically rely on static data inputs, often struggle to manage these time-dependent fluctuations effectively. This inefficiency can lead to decreased accuracy in analysis and misinterpretations of the user’s states, potentially resulting in services that do not meet user expectations.
III-C2 Energy Efficiency
Although the BCI-driven HMD alleviates the complexity of managing batteries for multiple sensors, processing real-time brain signals and constantly updating HDTs still requires significant computational power and energy resources. Current neural network architectures is energy-intensive, limiting their scalability and feasibility, particularly in large-scale, real-time applications within the Metaverse. Ensuring energy efficiency is critical to the sustainable operation of the Bio-HDT framework.
III-C3 Data Privacy and Security
Given the sensitive and personal nature of brain signals, ensuring their secure transmission and processing is crucial. Transmitting these signals directly over wireless networks presents significant challenges due to the susceptibility of these channels to noise and interference, which can greatly degrade the quality and integrity of the data.
Therefore, it is an urgent need to design efficient, secure, and scalable approaches for processing brain data in real-time. In the following section, we introduce an SNN-driven FL model designed to enhance Bio-HDT framework. This model offers a robust solution that not only enhances the efficacy of brain data processing but also minimizes energy consumption and bolsters privacy protections.
@@ IV Spiking Neural Network Enhanced Bio-HDT
Figure 3: The proposed network architecture based on SNNs. (a) Brain signal extraction is performed by HMD equipped with EEG electrodes to capture brain activity, (b) The captured brain signal is then input into a spike encoder, which converts the signal into spikes for further processing, (c) and (d) A brain signal identifier processes these firing spikes to accomplish brain signals identification.
@@ IV-A Spiking Neural Network
SNN is a bio-inspired neuromorphic computing model that closely mimics the way of human brain processes information. Different from traditional neural networks (i.e., ANNs) that process information continuously and require constant computational activity, SNNs leverage discrete spikes of electrical activity to transmit information between neurons, more accurately reflecting the communication process of biological neurons [11, 12].
In the context of the Bio-HDT framework and the human-centric Metaverse, SNNs offer significant advantages due to their ability to process information based on the timing and frequency of spikes, closely mirroring the brain’s natural information processing mechanisms. These capabilities enables SNNs particularly adept at interpreting rapidly fluctuating brain signals and recognizing dynamic, time-based patterns that traditional models often overlook. As illustrated in the step (2) of Fig. 2, extracted brain signals are first fed into a brain-like SNN model. This model learns to process and efficiently encode brain activity into multiple spikes, enhancing the real-time responsiveness and accuracy in user states analysis and personalized service delivery. Importantly, the event-driven nature of SNNs, i.e., activating only during specific spike events, significantly reduces energy consumption and computational overhead, enabling SNNs ideally suited for power-sensitive devices such as HMDs and sensors within the Bio-HDT framework (for Q2). More technical details of SNN can be found in [12, 13].
@@ IV-B SNN-Driven Federated Learning
MSPs traditionally collect users’ data to train deep learning models for states analysis. However, transmitting raw brain signals over wireless systems poses significant privacy concerns due to the sensitive and personal nature of this data. To address this challenge, we introduce the SNN-driven FL model. FL is a distributed machine learning approach where models are trained locally on user devices, and only the necessary updates are shared with a central server. This method eliminates the requirement to transmit raw data over the network, thereby enabling user privacy preservation while still enabling the collaborative training of models [14].
Specifically, SNN-driven FL can be deployed on local devices, i.e., BCI-driven HMDs. Instead of transmitting raw data to MSPs, each device trains its own SNN model using an FL framework. Note that the model training is conducted dynamically, leveraging idle periods or times of low interaction to process data and refine the model effectively. This strategy minimizes disruptions and maximizes efficiency during periods of active use. Once the local models are sufficiently trained, only the model updates (i.e., gradients) are transmitted to the MSP. There, they are aggregated to improve model performance without exposing the original brain data [14]. By retaining raw brain signals on local devices, the system protects users’ cognitive, emotional, and behavioral data from potential security threats during transmission. This architecture not only ensures a higher level of data privacy but also addresses critical concerns about the vulnerability of sensitive brain data (for Q3).

### v3

# Integrating Brain-Computer Interface and Neuromorphic Computing for Human Digital Twins

@@ 2. Q2)
    Given the need for continuous updating of HDT to adapt to real-time changes of user, what strategies can be developed to reduce the energy consumption involved in these data-intensive operations driven by deep learning model?
@@ 3. Q3)
  • We introduce a novel Bio-inspired HDT (Bio-HDT) framework that utilizes Brain-Computer Interface (BCI) technology to collect electroencephalography (EEG) signals from human brain as the primary data source for constructing HDTs. This framework harnesses EEG signals to capture real-time behavioral, psychological, physiological, and observational information of individuals, thereby providing richer data and a holistic view for constructing immersive and personalized HDTs. This approach not only enhances the precision and fidelity of the HDTs but also significantly reduces the reliance on heterogeneous IoT devices, simplifying both data collection and processing. Consequently, it ensures high QoS provided by MSPs and QoE perceived by users (For Q1).
  • We propose a Bio-inspired neuromorphic computing learning model based on the Spiking Neural Networks (SNNs) for efficiently processing extracted EEG signals. Unlike traditional Artificial Neural Networks (ANNs) that process information continuously and require constant computational activity, SNNs utilize discrete, brain-like spikes to transmit information between neurons. This method more accurately reflects the natural communication processes of brain neurons, thereby enhancing data processing efficiency and reducing the energy consumption specifically involved in the real-time updating of HDTs (For Q2).
@@ II-A Human-centric Metaverse and Human Digital Twin Framework
As the next generation of the Internet, the human-centric Metaverse extends beyond providing a virtual space for gaming. It is envisioned as a new paradigm that leverages users’ behavioral, psychological, physiological, and observational data to bridge the virtual and physical worlds, thereby profoundly transforming human lifestyles. HDTs, similar to the critical role of digital twins (DT) in manufacturing [rani2024human], play a vital role in improving the human-centric Metaverse. For instance, they allow simulation and prediction of physiological and psychological states of humans, facilitating personalized medical treatments [10238695] and health management programs [9839649].
Specifically, the current HDT framework operates cyclically by executing the following steps: (1) Sensors embedded in IoT devices continuously collect user real-time behavioral data; (2) The data are then transmitted to the MSP; (3) The MSP processes and analyzes the data using advanced software and algorithms. After analyzing these data, MSPs are able to generate the user’s HDT. Currently, this framework has been utilized primarily in healthcare. For example, HDTs are integrated with wearable devices such as smartwatches, fitness trackers, and health sensors to continuously monitor vital factors like heart rate, blood pressure, and glucose levels. This data can be synchronized with medical records and utilized to generate real-time insights into an individual’s health, enabling early diagnosis of potential issues and the optimization of personalized healthcare strategies [10238695, 9839649, asad2023human].
@@ II-B2 Energy Consumption (Q2)
To efficiently analyze these collected data, advanced models and algorithms based on deep learning have been deployed. However, they are primarily driven by ANNs, which are resource-intensive and require substantial computational power for both training and inference phases [wu2018development]. This intense energy demand presents significant challenges for these IoT devices with limited energy. First, the high power requirement for real-time data processing and analysis can lead to frequent battery depletion, disrupting continuous monitoring and data collection essential for maintaining accurate and dynamic HDTs. Moreover, the strain on device endurance due to persistent high-load operations can degrade device performance over time, increasing maintenance costs and potentially compromising the quality of user interaction within the Metaverse.
@@ III Brain-Computer Interface-Enhanced Human Digital Twin
In this section, we illustrate BCI technology and its current applications across various domains. We then propose a novel Bio-HDT framework that leverages advanced BCI technology. Finally, we discuss the unique challenges involved in deploying BCI within HDT frameworks.
@@ III-A Brain Signal and Brain-Computer Interface
Unlike previous works [10238695, 9839649, asad2023human], this work leverages brain signals as a novel data source for constructing HDTs. Specifically, brain signals offer a multifaceted view into human nature, crucial for modeling the human activities. They contain critical behavioral, psychological, physiological, and observational information, enabling the decoding of subconscious and conscious responses to environmental stimuli. These insights serve as windows into the internal processes that govern human actions, perceptions, and interactions, presenting their significant potential for constructing HDTs. By harnessing these signals, we can develop a human-centric Metaverse where each user’s HDT is authentically driven by their unique brain activities, ensuring a more immersive and personalized virtual experience.
To this end, BCI technology that connects the brain signals and physical devices has been considered. As a direct communication pathway between the human brain and externals machines, devices, and system, BCI operates by monitoring brain activities and collecting brain signals through sensors attached to the human scalp (e.g., non-invasive BCI [cognixion]), processing the data to detect specific patterns related to intended actions, and translating these patterns into actionable commands. This allows users to control devices or perform tasks without physical movement [neuralink_2024_neuralink]. It has been successfully applied in various domains, such as healthcare in arkinson’s [8119531] and Alzheimer’s [7967704], demonstrating its broad potential and utility. It has been successfully applied across multiple domains, including healthcare for Parkinson’s and Alzheimer’s disease [8119531, 7967704] and entertainment for thought-driven avatar control and immersive VR interaction [cognixion, 10496440], demonstrating broad potential and utility.
@@ III-B Biologically-inspired Human Digital Twins
Figure 1: The proposed framework for the biologically-inspired human digital twin (Bio-HDT).
@@
As analyzed above, brain signals encompass underlying information about human activities. These activities provide a holistic view of an individual’s real-time states across multiple dimensions, which is particularly useful in contexts such as healthcare, personalized technology interactions, and immersive Metaverse application. By capturing and analyzing these various human behavior and states information, BCI technology assists in creating more personalized and responsive systems that adapt to individual needs and changes in their condition. Therefore, we propose the Bio-HDT framework that utilizes brain signals as the data source for conducting HDT and incorporates BCI technology to efficiently capture these signals. In particular, this work primarily focuses on EEG signals due to their ease of non-invasive acquisition, high temporal resolution, and broad applicability in wearable BCI systems [10496440]. These characteristics make EEG an ideal modality for real-time monitoring and neural-digital interaction within the proposed Bio-HDT framework.
As illustrated in Fig. 1, the BCI-driven Head-mounted Display (HMD) system within the Bio-HDT framework continuously captures EEG signals from human brain through embedded non-invasive BCI sensors, which are then analyzed using a neuromorphic computing learning model that mimics brain function through SNNs. After real-time extraction and analysis of these EEG signals, the Bio-HDT is constructed and synchronized with the corresponding human states in the physical world. Utilizing this real-time Bio-HDT, MSPs can identify and even predict fluctuations in a user’s psychology and physical intentions. Consequently, this framework enables MSPs to generate personalized services for the users by monitoring the EEG signals acquired from their brain activity (for Q1).
@@
Specifically, the Bio-HDT operates as a two-layer bio-inspired scheme. The first layer bio-inspired captures EEG signals, which reflect the user’s cognitive, emotional, and behavioral states, providing an authentic representation of the brain’s biological processes. These EEG signals are continuously monitored and collected through embedded BCI sensors in the HMD system, ensuring real-time feedback from the user. The second layer bio-inspired processes these EEG signals using SNNs, a neuromorphic computing scheme that mimics the way neurons communicate in the brain. Different from traditional neural networks, SNNs process data in a manner that mirrors the brain’s natural signaling mechanisms, interpreting neural patterns through the timing and frequency of spikes, which is detailed in Section IV. This allows MSPs to accurately extract, analyze, and respond to users’ brain signals, generating personalized content in real time.
@@
By efficiently combining the BCI technology, the Bio-HDT framework offers several advantages. First, it minimizes the reliance on multiple IoT devices and external sensors traditionally required to monitor a user’s physical state. Conventional HDT systems rely on complex sensor networks, resulting in hardware demands, synchronization challenges, and increased energy consumption. By directly leveraging brain signals through the BCI, the Bio-HDT framework gathers critical multifaceted data (e.g., cognition, emotion, and behavior) without relying on an array of external sensors, thereby eliminating inefficiencies, reducing the investment in devices, and lowering computational load and data breach risk. Moreover, these EEG signals are extracted unconsciously, without requiring explicit physical actions, preserving the user’s immersive experience and providing a more nuanced representation of their state. This facilitates more natural and immersive interactions within the Metaverse.
@@
Figure 2: The proposed network architecture based on SNNs. (a) EEG signals extraction is performed by HMD equipped with EEG electrodes to capture brain activity, (b) The captured brain signal is then input into a spike encoder, which converts the signal into spikes for further processing, (c) and (d) A brain signal identifier processes these firing spikes to accomplish brain signals identification.
Therefore, the Bio-HDT framework not only addresses the challenge of customizing HDTs for individual users (for Q1) but also minimizes energy consumption (for Q2) and reduces the risk of data breaches (for Q3) by eliminating the need for multiple data collection devices. Additionally, the real-time interaction between brain signals and the decision-making systems enhances the accuracy, responsiveness, and immersiveness of HDTs. This new paradigm brings users closer to a fully realized human-centric Metaverse. It is worth noting that the proposed Bio-HDT system can also serve humans effectively beyond the Metaverse environment. In such cases, the Bio-HDT system can be used to monitor various aspects of human states [neuralink_2024_neuralink].
Although the BCI-based Bio-HDT offers rich signals for updating HDTs and advancing the human-centric Metaverse, practical deployment faces three bottlenecks: (i) decoding non-stationary brain activity, (ii) meeting tight energy budgets on always-on edge devices, and (iii) protecting sensitive neural data. These constraints call for efficient, secure, and scalable real-time brain-data pipelines. In the next section, we introduce an SNN-driven FL model that strengthens the Bio-HDT framework by improving dynamic-signal decoding, reducing energy via event-driven computation, and preserving privacy by keeping raw EEG on device.
@@ IV-A Spiking Neural Network
SNN is a bio-inspired neuromorphic computing model that closely mimics the way of human brain processes information. Different from traditional neural networks (i.e., ANNs) that process information continuously and require constant computational activity, SNNs leverage discrete spikes of electrical activity to transmit information between neurons, more accurately reflecting the communication process of biological neurons [9543525]. In the context of the Bio-HDT framework and the human-centric Metaverse, SNNs offer significant advantages due to their ability to process information based on the timing and frequency of spikes, closely mirroring the brain’s natural information processing mechanisms. These capabilities enables SNNs particularly adept at interpreting rapidly fluctuating brain signals and recognizing dynamic, time-based patterns that traditional models often overlook. As illustrated in the step (2) of Fig. 1, extracted brain signals are first fed into a brain-like SNN model. This model learns to process and efficiently encode brain activity into multiple spikes, enhancing the real-time responsiveness and accuracy in user states analysis and personalized service delivery. Importantly, the event-driven nature of SNNs, i.e., activating only during specific spike events, significantly reduces energy consumption and computational overhead, enabling SNNs ideally suited for power-sensitive devices such as HMDs and sensors within the Bio-HDT framework (for Q2). More technical details of SNN can be found in [shang2025energy].
In addition to efficiently processing brain signals, the flexibility of the proposed neuromorphic architecture allows it to be seamlessly extended to other data domains beyond BCI, including physiological, behavioral, and environmental modalities. By converting these diverse inputs into a unified spike-based representation, the SNN model can efficiently learn subtle temporal dynamics across heterogeneous data sources. This generalization enriches the Bio-HDT framework with more comprehensive human-state information and paves the way for multimodal HDT systems in future B5G/6G networks.
@@ IV-B SNN-Driven Federated Learning
MSPs traditionally collect users’ data to train deep learning models for states analysis. However, transmitting raw brain signals over wireless systems poses significant privacy concerns due to the sensitive and personal nature of this data. To address this challenge, we introduce the SNN-driven FL model. FL is a distributed machine learning approach where models are trained locally on user devices, and only the necessary updates are shared with a central server. This method eliminates the requirement to transmit raw data over the network, thereby enabling user privacy preservation while still enabling the collaborative training of models [McMahan2016CommunicationEfficient].
Specifically, SNN-driven FL can be deployed on local devices, i.e., BCI-driven HMDs. Instead of transmitting raw data to MSPs, each device trains its own SNN model using an FL framework. Once the local models are sufficiently trained, only the model updates (i.e., gradients) are transmitted to the MSP. At the MSP, the trained models are aggregated to improve model performance without exposing the original brain data [McMahan2016CommunicationEfficient]. By retaining raw brain signals on local devices, the system protects users’ cognitive, emotional, and behavioral data from potential security threats during transmission. This architecture not only ensures a higher level of data privacy but also addresses critical concerns about the vulnerability of sensitive brain data (for Q3).